OpenStudy (anonymous):

Use your calculator to determine the value of the following absolute value expression. Round your answer to the nearest hundredth. |2 - 2 i | ....I need help figuring out how to do this on a scientific calculator..

OpenStudy (anonymous):

2*Sqrt(2)

OpenStudy (anonymous):

or 2.83

OpenStudy (anonymous):

but what about the i?

OpenStudy (anonymous):

it = sqrt-1...

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^{2}+b^{2}}\]

OpenStudy (anonymous):

On a TI calculator it's abs(a+bi) I believe.

OpenStudy (anonymous):

So I would press abs then insert this equation for a+bi?

OpenStudy (anonymous):

I'm not seeing any abs key..

OpenStudy (anonymous):

it's in the math menu on mine and then the cpx menu I believe.

OpenStudy (anonymous):

I can't find it...but I have more math problems I need to get too...I need to take some time and play around with this online scientific calculator to figure it out probably..thanks anyway guys.

OpenStudy (anonymous):

Do you have a TI 84+ by any change?